# AI Agents vs. Agentic AI: A Conceptual Taxonomy, Applications and Challenges

### v1

@@ I Introduction
Building on this generative foundation, a new class of systems known as AI Agents emerged. These agents enhanced LLMs with capabilities for external tool use, function calling, and sequential reasoning, enabling them to retrieve real-time information and execute multi-step workflows autonomously [28, 29]. Frameworks such as AutoGPT [30] and BabyAGI (https://github.com/yoheinakajima/babyagi) exemplified this transition, showcasing how LLMs could be embedded within feedback loops to dynamically plan, act, and adapt in goal-driven environments [31, 32]. By late 2023, the field had advanced further into the realm of Agentic AI complex, multi-agent systems in which specialized agents collaboratively decompose goals, communicate, and coordinate toward shared objectives. Architectures such as CrewAI demonstrate how these agentic frameworks can orchestrate decision-making across distributed roles, facilitating intelligent behavior in high-stakes applications including autonomous robotics, logistics management, and adaptive decision-support [33, 34, 35, 36].
@@ I-A Methodology Overview
Building on these foundations, the review proceeds to the emergence of Agentic AI, which represents a significant conceptual leap. Here, we highlight the transformation from tool-augmented single-agent systems to collaborative, distributed ecosystems of interacting agents. This shift is driven by the need for systems capable of decomposing goals, assigning subtasks, coordinating outputs, and adapting dynamically to changing contexts capabilities that surpass what isolated AI Agents can offer.
@@ I-A1 Search Strategy
Targeted queries such as “Agentic AI + Coordination + Planning,” and “AI Agents + Tool Usage + Reasoning” were employed to retrieve papers addressing both conceptual underpinnings and system-level implementations. Literature inclusion was based on criteria such as novelty, empirical evaluation, architectural contribution, and citation impact. The rising global interest in these technologies illustrated in Figure 1 using Google Trends data reinforces the urgency of synthesizing this emerging knowledge space.
@@ II Foundational Understanding of AI Agents
AI Agents are an autonomous software entities engineered for goal-directed task execution within bounded digital environments [39, 14]. These agents are defined by their ability to perceive structured or unstructured inputs [40], reason over contextual information [41, 42], and initiate actions toward achieving specific objectives, often acting as surrogates for human users or subsystems [43]. Unlike conventional automation scripts, which follow deterministic workflows, AI agents demonstrate reactive intelligence and limited adaptability, allowing them to interpret dynamic inputs and reconfigure outputs accordingly [44]. Their adoption has been reported across a range of application domains, including customer service automation [45, 46], personal productivity assistance [47], internal information retrieval [48, 49], and decision support systems [50, 51].
@@ II-1 Overview of Core Characteristics of AI Agents
These core characteristics collectively enable AI Agents to serve as modular, lightweight interfaces between pretrained AI models and domain-specific utility pipelines. Their architectural simplicity and operational efficiency position them as key enablers of scalable automation across enterprise, consumer, and industrial settings. While limited in reasoning depth compared to more general AI systems, their high usability and performance within constrained task boundaries have made them foundational components in contemporary intelligent system design.
@@ II-2 Foundational Models: The Role of LLMs and LIMs
LLMs such as GPT-4 [73] and PaLM [74] are trained on massive datasets of text from books, web content, and dialogue corpora. These models exhibit emergent capabilities in natural language understanding, question answering, summarization, dialogue coherence, and even symbolic reasoning [75, 76]. Within AI agent architectures, LLMs serve as the primary decision-making engine, allowing the agent to parse user queries, plan multi-step solutions, and generate naturalistic responses. For instance, an AI customer support agent powered by GPT-4 can interpret customer complaints, query backend systems via tool integration, and respond in a contextually appropriate and emotionally aware manner [77].
@@
For example, as illustrated in Figure 5 in an autonomous drone agent tasked with inspecting orchards, a LIM can identify diseased fruits or damaged branches by interpreting live aerial imagery and triggering predefined intervention protocols. Upon detection, the system autonomously triggers predefined intervention protocols, such as notifying horticultural staff or marking the location for targeted treatment without requiring human intervention [57, 17]. This workflow exemplifies the autonomy and reactivity of AI agents in agricultural environment and recent literature underscores the growing sophistication of such drone-based AI agents. Chitra et al. [85] provide a comprehensive overview of AI algorithms foundational to embodied agents, highlighting the integration of computer vision, SLAM, reinforcement learning, and sensor fusion. These components collectively support real-time perception and adaptive navigation in dynamic environments. Kourav et al. [86] further emphasize the role of natural language processing and large language models in generating drone action plans from human-issued queries, demonstrating how LLMs support naturalistic interaction and mission planning. Similarly, Natarajan et al. [87] explore deep learning and reinforcement learning for scene understanding, spatial mapping, and multi-agent coordination in aerial robotics. These studies converge on the critical importance of AI-driven autonomy, perception, and decision-making in advancing drone-based agents.
@@ II-3 Generative AI as a Precursor
  • Prompt Dependency and Statelessness: Generative systems are stateless in that they do not retain context across interactions unless explicitly provided [98, 99]. Their design lacks intrinsic feedback loops [100], state management [101, 102], or multi-step planning a requirement for autonomous decision-making and iterative goal refinement [103, 104].
@@ II-A2 Tool-Augmented AI Agents: Enhancing Functionality
Tool Invocation. When an agent identifies a need that cannot be addressed through its internal knowledge such as querying a current stock price, retrieving up-to-date weather information, or executing a script, it generates a structured function call or API request [124, 125]. These calls are typically formatted in JSON, SQL, or Python, depending on the target service, and routed through an orchestration layer that executes the task.
Result Integration. Once a response is received from the tool, the output is parsed and reincorporated into the LLM’s context window. This enables the agent to synthesize new reasoning paths, update its task status, and decide on the next step. The ReAct framework [126] exemplifies this architecture by combining reasoning (Chain-of-Thought prompting) and action (tool use), with LLMs alternating between internal cognition and external environment interaction.
@@ V Challenges and Limitations in AI Agents and Agentic AI
Figure 12: Challenges and Solutions Across Agentic Paradigms. (a) Key limitations of AI Agents including causality deficits and shallow reasoning. (b) Amplified coordination and stability challenges in Agentic AI systems.

### v2

@@ I Introduction
Although the term AI agent was first introduced in 1998 [3], it has since evolved significantly with the rise of generative AI. Building upon this generative foundation, a new class of systems—commonly referred to as AI agents—has emerged. These agents enhanced LLMs with capabilities for external tool use, function calling, and sequential reasoning, enabling them to retrieve real-time information and execute multi-step workflows autonomously [28, 29]. Frameworks such as AutoGPT [30] and BabyAGI (https://github.com/yoheinakajima/babyagi) exemplified this transition, showcasing how LLMs could be embedded within feedback loops to dynamically plan, act, and adapt in goal-driven environments [31, 32]. By late 2023, the field had advanced further into the realm of Agentic AI complex, multi-agent systems in which specialized agents collaboratively decompose goals, communicate, and coordinate toward shared objectives. In line with this evolution, Google introduced the Agent-to-Agent (A2A) protocol in 2025 [33], a proposed standard designed to enable seamless interoperability among agents across different frameworks and vendors. The protocol is built around five core principles: embracing agentic capabilities, building on existing standards, securing interactions by default, supporting long-running tasks, and ensuring modality agnosticism. These guidelines aim to lay the groundwork for a responsive, scalable agentic infrastructure.
Architectures such as CrewAI demonstrate how these agentic frameworks can orchestrate decision-making across distributed roles, facilitating intelligent behavior in high-stakes applications including autonomous robotics, logistics management, and adaptive decision-support [34, 35, 36, 37].
@@ I-A Methodology Overview
Building on these foundations, the review proceeds to the emergence of Agentic AI, which represents a significant conceptual leap. Here, we highlight the transformation from tool-augmented single-agent systems to collaborative, distributed ecosystems of interacting agents. This shift is driven by the need for systems capable of decomposing goals, assigning subtasks, coordinating outputs, and adapting dynamically to changing contexts—capabilities that surpass what isolated AI Agents can offer.
@@ I-A1 Search Strategy
Targeted queries such as “Agentic AI + Coordination + Planning,” and “AI Agents + Tool Usage + Reasoning” were employed to retrieve papers addressing both conceptual underpinnings and system-level implementations. Literature inclusion was based on criteria such as novelty, empirical evaluation, architectural contribution, and citation impact. The rising global interest in these technologies, as illustrated in Figure 1 using Google Trends data, underscores the urgency of synthesizing this emerging knowledge space.
@@ II Foundational Understanding of AI Agents
AI Agents are an autonomous software entities engineered for goal-directed task execution within bounded digital environments [40, 14]. These agents are defined by their ability to perceive structured or unstructured inputs [41], reason over contextual information [42, 43], and initiate actions toward achieving specific objectives, often acting as surrogates for human users or subsystems [44]. Unlike conventional automation scripts, which follow deterministic workflows, AI agents demonstrate reactive intelligence and limited adaptability, allowing them to interpret dynamic inputs and reconfigure outputs accordingly [45]. Their adoption has been reported across a range of application domains, including customer service automation [46, 47], personal productivity assistance [48], internal information retrieval [49, 50], and decision support systems [51, 52]. A noteworthy example of autonomous AI agents is Anthropic’s ”Computer Use” project, where Claude was trained to navigate computers to automate repetitive processes, build and test software, and perform open-ended tasks such as research [53].
@@ II-1 Overview of Core Characteristics of AI Agents
These core characteristics collectively enable AI Agents to serve as modular, lightweight interfaces between pretrained AI models and domain-specific utility pipelines. Their architectural simplicity and operational efficiency position them as key enablers of scalable automation across enterprise, consumer, and industrial settings. Although still limited in reasoning depth compared to more general AI systems [75], their high usability and performance within constrained task boundaries have made them foundational components in contemporary intelligent system design.
@@ II-2 Foundational Models: The Role of LLMs and LIMs
LLMs such as GPT-4 [76] and PaLM [77] are trained on massive datasets of text from books, web content, and dialogue corpora. These models exhibit emergent capabilities in natural language understanding, question answering, summarization, dialogue coherence, and even symbolic reasoning [78, 79]. Within AI agent architectures, LLMs serve as the primary decision-making engine, allowing the agent to parse user queries, plan multi-step solutions, and generate naturalistic responses. For instance, an AI customer support agent powered by GPT-4 can interpret customer complaints, query backend systems via tool integration, and respond in a contextually appropriate and emotionally aware manner [80, 81].
@@
For example, as illustrated in Figure 5 in an autonomous drone agent tasked with inspecting orchards, a LIM can identify diseased fruits [89] or damaged branches by interpreting live aerial imagery and triggering predefined intervention protocols. Upon detection, the system autonomously triggers predefined intervention protocols, such as notifying horticultural staff or marking the location for targeted treatment without requiring human intervention [59, 17]. This workflow exemplifies the autonomy and reactivity of AI agents in agricultural environment and recent literature underscores the growing sophistication of such drone-based AI agents. Chitra et al. [90] provide a comprehensive overview of AI algorithms foundational to embodied agents, highlighting the integration of computer vision, SLAM, reinforcement learning, and sensor fusion. These components collectively support real-time perception and adaptive navigation in dynamic environments. Kourav et al. [91] further emphasize the role of natural language processing and large language models in generating drone action plans from human-issued queries, demonstrating how LLMs support naturalistic interaction and mission planning. Similarly, Natarajan et al. [92] explore deep learning and reinforcement learning for scene understanding, spatial mapping, and multi-agent coordination in aerial robotics. These studies converge on the critical importance of AI-driven autonomy, perception, and decision-making in advancing drone-based agents.
@@ II-3 Generative AI as a Precursor
  • Prompt Dependency and Statelessness: Although generative systems are stateless in that they do not retain context across interactions unless explicitly provided [103, 104], recent advancements like GPT-4.1 support larger context windows-up to 1 million tokens-and are better able to utilize that context thanks to improved long-text comprehension [105]. Their design also lacks intrinsic feedback loops [106], state management [107, 108], or multi-step planning a requirement for autonomous decision-making and iterative goal refinement [109, 110].
@@ II-A2 Tool-Augmented AI Agents: Enhancing Functionality
Tool Invocation. When an agent identifies a need that cannot be addressed through its internal knowledge such as querying a current stock price, retrieving up-to-date weather information, or executing a script, it generates a structured function call or API request [130, 131]. These calls are typically formatted in JSON, SQL, or Python dictionary, depending on the target service, and routed through an orchestration layer that executes the task.
Result Integration. Once a response is received from the tool, the output is parsed and reincorporated into the LLM’s context window. This enables the agent to synthesize new reasoning paths, update its task status, and decide on the next step. The ReAct framework [132] exemplifies this architecture by combining reasoning (Chain-of-Thought prompting) and action (tool use), with LLMs alternating between internal cognition and external environment interaction. A prominent example of a tool-augmented AI agent is ChatGPT, which, when unable to answer a query directly, autonomously invokes the Web Search API to retrieve more recent and relevant information, performs reasoning over the retrieved content, and formulates a response based on its understanding [133].
@@ V Challenges and Limitations in AI Agents and Agentic AI
Figure 12: Illustration of Chellenges: (a) Key limitations of AI Agents including causality deficits and shallow reasoning. (b) Amplified coordination and stability challenges in Agentic AI systems.
@@ VI Potential Solutions and Future Roadmap
A transformative direction for future AI systems is introduced by the Absolute Zero: Reinforced Self-play Reasoning with Zero Data (AZR) framework, which reimagines the learning paradigm for AI Agents and Agentic AI by removing dependency on external datasets [216]. Traditionally, both AI Agents and Agentic AI architectures have relied on human-annotated data, static knowledge corpora, or preconfigured environments factors that constrain scalability and adaptability in open-world contexts. AZR addresses this limitation by enabling agents to autonomously generate, validate, and solve their own tasks, using verifiable feedback mechanisms (e.g., code execution) to ground learning. This self-evolving mechanism opens the door to truly autonomous reasoning agents capable of self-directed learning and adaptation in dynamic, data-scarce environments.
In the context of Agentic AI—where multiple specialized agents collaborate within orchestrated workflows AZR lays the groundwork for agents to not only specialize but also co-evolve. For instance, scientific research pipelines could consist of agents that propose hypotheses, run simulations, validate findings, and revise strategies—entirely through self-play and verifiable reasoning, without continuous human oversight. By integrating the AZR paradigm, such systems can maintain persistent growth, knowledge refinement, and task flexibility across time. Ultimately, AZR highlights a future in which AI agents transition from static, pretrained tools to intelligent, self-improving ecosystems—positioning both AI Agents and Agentic AI at the forefront of next-generation artificial intelligence.